molal depression of freezing point of water is 1.86 per 1000g of water 0.02 mol of urea dissolved in 100g of water will produce a lowering of temperature of
 

Grade:8

1 Answers

Vikas TU
4 years ago
Dear student 
 
Del Tf = (1000*Kf * w)/ (m*w)
Del Tf = ( 1000* 1.86 * 342 )/(1000*342)
=> 1.86 
Tf = 0- 1.86 = -1.86 
Hope this helps 
Cheers 
Good luck